The masses of articles produced in a particular workshop are normally distributed with mean A and standard deviation B. 5% of the article have a mass greater than 85g and 10% have a mas less than 25g. Find the values of A and B.
1
Expert's answer
2020-06-11T19:54:35-0400

Let X = the random variable denoting the mass of an article


Then by the problem,


X ~ N(A, B2)


Therefore, we know,


Z = "\frac{X-A}{B}" ~ N(0, 1)


Now, 5% of the article have a mass greater than 85g


i.e. the probability that an article having mass greater than 85g is 5%


i.e. P(X > 85) = 5%


i.e. P("\frac{X-A}{B}>\frac{85-A}{B}") = 0.05


i.e. 1 - P("\frac{X-A}{B}\leq\frac{85-A}{B}") = 0.05


i.e. P("Z\leq\frac{85-A}{B}") = 1 - 0.05 = 0.95


i.e. "\Phi(\frac{85-A}{B})" = 0.95 = "\Phi(1.64)"


i.e. "\frac{85-A}{B}=1.64"


i.e. A + 1.64B = 85 .......................(1)


Again, 10% of the article have a mas less than 25g


i.e. the probability that an article having mass less than 25g is 10%


i.e. P(X < 25) = 10%


i.e. P("\frac{X-A}{B}" <"\frac{25-A}{B}"​) = 0.1


i.e. P(Z <"\frac{25-A}{B}"​) = 0.1


i.e. "\Phi(\frac{25-A}{B})" = 0.1 = "\Phi(-1.28)"


i.e. "\frac{25-A}{B}" = - 1.28


i.e. A - 1.28B = 25 ........................(2)


Solving (1) and (2) we get, A = 51.3, B = 20.55


Answer: The values of A and B are 51.3g and 20.55g respectively.